Let X1,X2,.....,Xbe random sample of size n from a distribution with probability 

density function 

f(X; ,0)={ θX^(θ-1), 0<X<1, θ>0; 0, elsewhere

Obtain a maximum likeyhood estimator of θ.


1
Expert's answer
2021-06-04T11:21:15-0400

Solution:

Given,

"f(x; θ)=\{ θx^{θ-1}, 0<x<1, θ>0
\\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 0, \ elsewhere"

"L(θ)= \Pi_{i=1}^n \ f(x_i; θ)= \Pi_{i=1}^n \ θx_i^{θ-1}
\\=θx_1^{θ-1}.θx_2^{θ-1}.θx_3^{θ-1}....θx_n^{θ-1}
\\=θ^n{x_i}^{θ-1}"

"\Rightarrow \log L(θ)=\log (θ^n{x_i}^{θ-1})=\log (θ^n)+\log ({x_i}^{θ-1})
\\=n \logθ+(θ-1)\log x_i"

"\Rightarrow \dfrac{d \ \log L(θ)}{dθ}=\dfrac nθ+\log x_i"

Put "\dfrac{d \ \log L(θ)}{dθ}=0"

"\dfrac nθ+\log x_i=0
\\ \Rightarrow \dfrac nθ=-\log x_i
\\ \Rightarrow θ=\dfrac{n}{-\log x_i}=\dfrac{-n}{\log (\dfrac{x_i}{n}.n)}=\dfrac{-n}{(\log \bar x)+(\log n)}"

Thus, MLE "=\hat θ=\dfrac{-n}{(\log \bar x)+(\log n)}"